OpenStudy (anonymous):

solve the system algebraically x^2+y^2=36 x^2-y^2=9

OpenStudy (anonymous):

add those two equations to get: 2x^2 = 45 these are your x-coordinates of where the intersections occur. then plug them back in to get your y-coordinates.

jimthompson5910 (jim_thompson5910):

Add the equations: x^2+y^2=36 x^2-y^2=9 ------------ 2x^2 + 0y^2 = 45 So we have 2x^2 = 45, solve this for x 2x^2 = 45 x^2 = 45/2 x = sqrt(45/2) or x = -sqrt(45/2) x = sqrt(45)/sqrt(2) or x = -sqrt(45)/sqrt(2) x = (sqrt(45)*sqrt(2))/sqrt(2)*sqrt(2)) or x = -(sqrt(45)*sqrt(2))/sqrt(2)*sqrt(2)) x = sqrt(45*2)/sqrt(2*2) or x = -sqrt(45*2)/sqrt(2*2) x = sqrt(90)/sqrt(4) or x = -sqrt(90)/sqrt(4) x = sqrt(3^2*10)/sqrt(2^2) or x = -sqrt(3^2*10)/sqrt(2^2) x = sqrt(3^2)*sqrt(10)/sqrt(2^2) or x = -sqrt(3^2)*sqrt(10)/sqrt(2^2) x = 3*sqrt(10)/2 or x = -3*sqrt(10)/2 Now that you know the value of x, you can use it to find the value of y. So plug in x = 3*sqrt(10)/2 into either equation to find the corresponding value for y Also, plug in x = -3*sqrt(10)/2 to find the other solution for y.

OpenStudy (anonymous):

i think elimination is meant to be used, thats more graphically

OpenStudy (anonymous):

side note: the system is that of a circle and hyperbola both centered at the origin. they intersect at 4 points so that should coincide with the number of solutions you get.

OpenStudy (anonymous):

what u talking about? when you add those equations, you eliminated the y^2... i did use elimination...

OpenStudy (anonymous):

x=22.5 , y= 21.68